OpenStudy (anonymous):

A sphere has a diameter of 18 meters. Find the volume of the sphere, in cubic meters, in terms of π.

OpenStudy (anonymous):

\[4/3\Pi(r^{3})\] is the volume of the sphere. The diameter is 18, so the radius is 9. \[4/3\Pi(9^{3})\] = \[4/3\Pi(729})\] = \[4/3(2290)\] = 3053.63

OpenStudy (anonymous):

thank you sooo muchh :D

OpenStudy (anonymous):

wait is that it in the box?

OpenStudy (anonymous):

You got it. The key is to remember that the diameter has to be cut in half. The answer is 3053.63... just trying to get used to the coding here. That's (4/3)*PI(729)

OpenStudy (anonymous):

okay but in terms of pi is the one in the box right? and the other is simplified?

OpenStudy (anonymous):

Correct.